OpenStudy (anonymous):

find the value of cot(11pi/6)

OpenStudy (jlg030597):

the exact value or just value

OpenStudy (anonymous):

my paper only says value so thats what im going off of

OpenStudy (jlg030597):

remember that 11pi/6 is in 4th quadrant, so only cos and sec ratio will be positive... cot(11pi/6)=-cot (2pi-11pi/6)---->as cot x=-cot(-x)=-cot (2pi-x) =-cot(pi/6)=- sqrt(3) now cos(11pi/6)=cos (2pi-11pi/6)---->as cos x=cos(-x)=cos (2pi-x) =cos(pi/6)=sqrt(3)/2 csc(11pi/6)=-csc (2pi-11pi/6)---->as csc x=-csc(-x)=-csc (2pi-x) =-csc(pi/6)=- 2

OpenStudy (anonymous):

i know the answer is the sqrt of 3 i just dont get were the -cos and the 2 comes from

OpenStudy (jlg030597):

Well is this for Geometry.

OpenStudy (anonymous):

advance function and modeling

OpenStudy (jlg030597):

Oh then sorry thats alittle more advanced than i have gone with math.

OpenStudy (anonymous):

ok well thanks anyways

OpenStudy (jlg030597):

Yah Im only in Geometry and the highest math ive worked with is precal and that was for an engineering class.